OpenStudy (anonymous):

Ax+By=C what is the result if A=0

jigglypuff314 (jigglypuff314):

umm if you plug in 0 for A then (0)x + By = C so By = C then divide both sides by B so that y = C/B

OpenStudy (anonymous):

Thankss !

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

What it'd be if B=0?

jigglypuff314 (jigglypuff314):

same idea :) if you plug in 0 for B then Ax + (0)y = C so Ax = C then divide both sides by A so that y = C/A

OpenStudy (anonymous):

Oh I knew that...... :\ what kind of graph would it be tho ?

jigglypuff314 (jigglypuff314):

if A = 0 it would be y = something so it would be a horizontal if B = 0 it would be x = something so it would be vertical

OpenStudy (anonymous):

OMG Thankss ! :D

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

Do you know this if the x-intercept of the graph of 2x+6y=4 is 2 is that true or false ?

jigglypuff314 (jigglypuff314):

x intercept is when y = 0 so when you plug in 0 for y 2x + 6(0) = 4 so 2x = 4 then divide both sides by 2 so x = 2

OpenStudy (anonymous):

Thanks !! sorry if i ask to much ! but do you know .. If the y-intercept of the graph of x-2y=7 is 3.5

jigglypuff314 (jigglypuff314):

y intercept is when x = 0 so plug in 0 for x so (0) - 2y = 7 -2y = 7 then divide both sides by -2 so y = -7/2 = negative 3.5

OpenStudy (anonymous):

Thankss !!! Can you help me on how to find slope ?...

jigglypuff314 (jigglypuff314):

sure :) rise ( y2 - y1 ) slope = ----- = ---------- run ( x2 - x1 ) when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

Yeah like i have the point (-2,5), (8,-4)

jigglypuff314 (jigglypuff314):

ok then rise ( y2 - y1 ) ( (-4) - (5) ) ( -9 ) slope = ----- = ---------- = ----------- = ----- run ( x2 - x1 ) ( (8) - (-2) ) ( 10 )

jigglypuff314 (jigglypuff314):

if you're wondering where the numbers came from: (x1, y1) and (x2, y2) (-2 , 5) and (8 , -4) ^ ^ ^ ^ that's how I matched them up :)

OpenStudy (anonymous):

Thanks !! what if the points are (6,-2), (-3,7)

jigglypuff314 (jigglypuff314):

(x1, y1) and (x2, y2) (6 , -2) and (-3 , 7) so rise ( y2 - y1 ) ( (7) - (-2) ) ( 9 ) slope = ----- = ---------- = ----------- = ----- = -1 run ( x2 - x1 ) ( (-3) - (6) ) ( -9 )

OpenStudy (anonymous):

Thanks ! how would you write (10,1); m=1/5 in point slope form ?

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when given m = slope and point (x1, y1) so plug in m = 1/5 and x1=10 and y1=1 into y - (y1) = (m) (x - (x1) )

OpenStudy (anonymous):

so it'd be y-(10)=(1/5)(x-1) right ?

jigglypuff314 (jigglypuff314):

close :) y - (1) = (1/5) (x - 10)

OpenStudy (anonymous):

Oh i see I read it wrong lmao

jigglypuff314 (jigglypuff314):

yep ^.^

OpenStudy (anonymous):

Yeah how would you write an equation in slope intercept form if the points are (-2,2) and (2,-8)

jigglypuff314 (jigglypuff314):

well first find the slope :) give it a try

OpenStudy (anonymous):

wait how do we find the slope again i forgot ...

jigglypuff314 (jigglypuff314):

rise ( y2 - y1 ) slope = ----- = ---------- = run ( x2 - x1 )

jigglypuff314 (jigglypuff314):

when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

oh yeahh thanks !!

OpenStudy (anonymous):

how do you know which one is x1 and x2 and y1 and y2

jigglypuff314 (jigglypuff314):

whatever way it lines up :P (x1, y1) and (x2, y2) (-2, 2 ) and ( 2, -8)

OpenStudy (anonymous):

Ohhh i feel dumb lmao

jigglypuff314 (jigglypuff314):

lol it's fine :) just be careful to not mess up that something like 2 - (-2) = 2+2

OpenStudy (anonymous):

the slope is -10/4 right ....?

jigglypuff314 (jigglypuff314):

yep! :) which can be simplified to -2.5 then slope intercept form is y = mx + b where m = slope so when m = -2.5 y = -2.5x + b but then we need to find b so we temporarily plug in a given point like (-2, 2) so 2 = (-2.5)(-2) + b solve for b

OpenStudy (anonymous):

how would we solve for b?

jigglypuff314 (jigglypuff314):

2 = (-2.5)(-2) + b 2 = 5 + b then subtract 5 from both sides to get b

OpenStudy (anonymous):

so -3=b

jigglypuff314 (jigglypuff314):

yep :) which we plug back into y = -2.5x + b to get y = -2.5x - 3 tah-dah! :)

OpenStudy (anonymous):

Omg Thankss !! you're like a life saver !!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

do you know how to find x and y intercepts ?

jigglypuff314 (jigglypuff314):

x intercept is when y = 0 y intercept is when x = 0

OpenStudy (anonymous):

oh wait so how'd you find 4x-3y=12 ?

jigglypuff314 (jigglypuff314):

x intercept is when y = 0 so 4x - 3(0) = 12 4x = 12 x = 3 is the x intercept then y intercept is when x = 0 so 4(0) - 3y = 12 -3y = 12 y = -4 is the y intercept

OpenStudy (anonymous):

Oh okay thanks ! how would you like graph an equation if it's like y=-3/4x+2

jigglypuff314 (jigglypuff314):

y intercept would be y = 2 in other words at point (0, 2) and x intercept would be at x = 8/3 in other words at point (8/3, 0) you would plot those points on the graph then draw a line through it :)

OpenStudy (anonymous):

Oh okay what if it's 2x+3y=9

jigglypuff314 (jigglypuff314):

y intercept would be y = 3 in other words at point (0, 3) x intercept would be x = 9/2 in other words at point (9/2, 0) you would plot those points on the graph then draw a line through it :)

OpenStudy (anonymous):

THANK YOUUU !!!!

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

what if we're doing like slope intercept form but it doesn't show the points like this one x-4y=8

jigglypuff314 (jigglypuff314):

slope intercept form is y = mx + b so in a sense, it just wants you to get "y" alone then :) so x - 4y = 8 -4y = -x + 8 y = (1/4)x - 2

OpenStudy (anonymous):

Oh how about y-2=5(x-1)

jigglypuff314 (jigglypuff314):

same idea :) get y alone and simplify

OpenStudy (anonymous):

Oh! so it'd be y=5x-1+2 Right...?

jigglypuff314 (jigglypuff314):

remember to distribute the 5 :)

OpenStudy (anonymous):

Oh ! so it'd be y=5x-5+2

jigglypuff314 (jigglypuff314):

then add like terms :)

OpenStudy (anonymous):

? how would you do that ..? lmao

jigglypuff314 (jigglypuff314):

well you can add -5 + 2 right? :P

OpenStudy (anonymous):

I guess lol

jigglypuff314 (jigglypuff314):

so you'll get y = 5x - 3

OpenStudy (anonymous):

Oh lmao i'm bad at math as you can see lol

jigglypuff314 (jigglypuff314):

lol you're doing really well compared to some other people I've had to deal with XD trust me :P

OpenStudy (anonymous):

how'd you write y=-2x+3 in standard form...? lol sorry if i ask a lot

OpenStudy (anonymous):

and lmao really i feel smart and dumb lmao but thanks :)

jigglypuff314 (jigglypuff314):

standard form is ax + by = c so basically get all the values with variables attached to them to one side so for y = -2x + 3 just add 2x to both sides :)

OpenStudy (anonymous):

Ohh so it'd be 2x+y=3

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

wait so if i do y-4=2(x-1) we distribute the 2 which will now be y-4=2x-2right ? and after that what do we do?

jigglypuff314 (jigglypuff314):

yep :) then add 4 to both sides then subtract 2x from both sides :)

OpenStudy (anonymous):

y=2x-2+4 right ?

jigglypuff314 (jigglypuff314):

yeah then subtract 2x from both sides :)

OpenStudy (anonymous):

y-2x=2+4 right

OpenStudy (anonymous):

then we subtract 4 ?...

jigglypuff314 (jigglypuff314):

mmm -2x + y = 2 2x - y = -2

OpenStudy (anonymous):

Thank Youu !!! :D i felt like you're a better teacher than my math teacher lol like no joke

jigglypuff314 (jigglypuff314):

thanks :) lol

OpenStudy (anonymous):

how'd we graph y=lxl-1 ....?

OpenStudy (anonymous):

and lmaoo your welcome :)

jigglypuff314 (jigglypuff314):

original graph was https://lh3.ggpht.com/_F6M745sbI3g/TI-Re7FDUwI/AAAAAAAAABM/6Q-oNeWjmxI/s1600/absolute+value.gif then y = |x| - 1 is that graph shifted down 1

OpenStudy (anonymous):

Oh okay Thankss !!

jigglypuff314 (jigglypuff314):

no problem :)